two identical particles carry a charge Q each. initially one is at rest on a smooth horizontal and the other is projected along the plane directly towards the first particle from a large distance with a speed v. what is the closest distance of approach ?

Grade:

2 Answers

Askiitians Expert Soumyajit IIT-Kharagpur
13 years ago

Dear Anand Shukla,

Ans:- As there is no external force present hence the system's total energy will remain constant. So, the initial KE will be totally converted to the system's PE when they are closest to each other. Hence from there we get,

(1/4∏ε)Q²/R=(1/2)mv²

R=Q²/(2∏εv²)

 

 

Please feel free to post as many doubts on our discussion forum as you can. If you find any question
Difficult to understand - post it here and we will get you the answer and detailed solution very quickly. We
are all IITians and here to help you in your IIT JEE preparation.


All the best Anand Shukla !!!

 


Regards,

Askiitians Experts
Soumyajit Das IIT Kharagpur

Kushagra Madhukar
3 years ago
Dear student,
Please find the solution to your problem.
 
As there is no external force present hence the system's total energy will remain constant. So, the initial KE will be totally converted to the system's PE when they are closest to each other. Hence from there we get,
(1/4πε)Q²/R
= (1/2)mv²
R = Q²/(2πεv²)
 
Thanks and regards,
Kushagra